calculate the tension in cable and reaction forces at pin

For the structure loaded as shown below, calculate
Answer #1

The net torque about point c is zero

torque T = F * perpendicular distance

or T = r * F * sin(theta)

the net torque about point c is

-(TAB * sin(25) * 4L) + (3200*3L) + (900*2L) = 0

TAB = 6743.67 lb

2)

The horizontal component of tension is

cx = TAB * cos(25)

cx = 6111.84 lb

3)

the vertical component of tension is

Ty = TAB * sin(25)

Ty = 2850 lb

the vertical reaction force at c is

cy = 3200 + 900 - Ty

cy = 4100 - 2850

cy = 1250 lb

25°.1